# From tug-of-war to Brownian Boost: explicit ODE solutions for player-funded stochastic-differential games

### v3

@@ 1.6. Brownian Boost and the high-noise limit
We now return to Brownian Boost. Our analysis of the game operates by regularizing it via a fine-mesh high-noise scaling of the Trail of Lost Pennies. In a first subsection, we advocate this discretization as a natural means of rigorous analysis of Brownian Boost. In the second, we complete the presentation of our principal conclusions about $\mathrm{BB}(\rho)$ by stating Theorem 1.18, which describes equilibrium stakes and gameplay in the high-noise regime via the analytic framework of Theorem 1.3 and Definition 1.4.
@@ 1.6.1. The scaled high-noise trail game as a regularized Brownian Boost
As such, $\mathrm{BB}_{\kappa}(\rho)$ may be more closely compared to a scaled version $\mathrm{ScTLP}(\kappa,\rho)$ of the Trail of Lost Pennies. The scaled game operates by the rules of $\mathrm{TLP}(\kappa,\rho)$ with two changes: it is played on $\kappa\,\mathbb{Z}$, not $\mathbb{Z}$; and the running costs (3) that enter the net receipt formulas (4) now include $\kappa^{2}$-prefactors, $C_{\pm}=\kappa^{2}\sum_{k=0}^{\infty}S_{\pm}(k)$.
@@
It is natural to pose the problem of determining the ‘domain of attraction’ of discretized approximant games for $\mathrm{BB}(\rho)$. Adapting the methods of [13, 15], [20] addresses this type of question for stochastic differential games of zero-sum; while the strongly non-anticipatory framework in [9] is adapted to the non-zero-sum case. Implementing a framework such as [9]’s rigorously for $\mathrm{BB}(\rho)$ would require careful handling of the infinite horizon, non-zero-sum payoffs, and the absence of discounting in Brownian Boost. Instead we choose the concrete discretization $\mathrm{TLP}(\kappa,\rho)$ in the limit of low $\kappa$ as the rigorous point of contact with $\mathrm{BB}(\rho)$.
@@ 1.6.2. Scaled gameplay in the high-noise trail game
Here we substantiate that for $\rho\in(0,1]$ the time-homogeneous Markov-perfect equilibria of $\mathrm{BB}(\rho)$ are given by the prescription in Definitions 1.2 and 1.4, with a result showing that this description captures (in the limit of low $\kappa$) all time-invariant stake-profiles and gameplay in the scaled trail game $\mathrm{ScTLP}(\kappa,\rho)$.
@@
is the unique element of ABMN($\kappa,\rho$) with $\phi_{0}=x$ and $m_{-1,0}=\kappa$ (as well as $m_{-\infty}=n_{\infty}=0$). In view of Theorems 1.8 and  1.16, time-invariant Nash equilibria in $\mathrm{TLP}(\kappa,\rho)$ are characterized up to the trivial symmetries by these solutions, and we use them to express our result Theorem 1.18.
@@
    $\kappa^{-2}a^{\rm def}_{\lfloor\kappa^{-1}r\rfloor}(\kappa,\rho,x)$ $\,=\,a_{\rho}(x,r)\;\big(1+\ O(\kappa)\big)\,$, (14)
    $\kappa^{-2}b^{\rm def}_{\lfloor\kappa^{-1}r\rfloor}(\kappa,\rho,x)$ $\,=\,b_{\rho}(x,r)\;\big(1+\ O(\kappa)\big)\,$, (15)
    with the implicit constant in the $O$-terms being uniform in $(\rho,x,r)\in(0,1]\times K\times K$ for compact $K\subset(0,\infty)$.
@@
    $0,\infty)\to\mathbb{R}:u\to\kappa\,X_{\kappa,\rho}(\kappa^{-2}z,\kappa^{-2}u)\,$,
    whose domain of definition is enlarged from $\kappa\,\mathbb{N}$ to $0,\infty)$ by interpolation.
    Equip the space $\mathcal{C}$ of continuous functions $f:[0,\infty)\to\mathbb{R}$ with the topology of uniform convergence on compact intervals. As $\kappa\searrow 0$, this process converges weakly in $\mathcal{C}$ to the unique solution $Z_{u}$ of the stochastic differential equation
    ${\rm d}Z_{u}=R_{\rho}(u)\,{\rm d}u+{\rm d}W_{u},$
    with $Z_{0}=z$, where $W_{u}$ is standard Brownian motion. The drift coefficient $R_{\rho}(u)$ equals $\frac{1-S_{\rho}(1,u)^{\rho}}{1+S_{\rho}(1,u)^{\rho}}$ or equivalently $\frac{1-S_{1}(1,\rho^{2}u)}{1+S_{1}(1,\rho^{2}u)}$. It has asymptotics
@@
Remark. The function $J(u)$ in the earlier signpost (1) equals $S_{1}(1,\rho^{2}u)$. The form for $J^{\prime}$ recorded there is given by taking $\rho=1$ in Definition 1.2 with a linear change of variable: see Lemma 5.1.
@@ 1.7.1. Fixed-parameter ABMN($\kappa,\rho$) asymptotics, and asymmetric decay
These themes are apparent in $\mathrm{BB}(\rho)$ from Proposition 1.5, wherein the choice $x=1$ locates the battlefield region in which stake-profiles are $\Theta(1)$ in a compact neighbourhood of the origin, while satisfying
$a_{-u}\ll b_{-u}\ll a_{-u}/b_{-u}\ll 1\,\,\,\,\textrm{for}\,\,\,\,u\gg 0:$ (16)
in negative territory, where Mina leads, stakes have fallen exponentially, the more so for Maxine, though the decay in stake ratio has a more modest polynomial rate. ‘Battlefield Cyl Fog’ (cut your losses, foot on gas) is a mnemonic for the premise (16), the phrases descriptive of the trailing and leading player’s respective approach far from the battlefield.
In [26], TLP(1,1)(1,1) was studied: the Trail of Lost Pennies without flip moves whose turn outcomes are decided by the simple $\frac{a}{a+b}$ lottery rule. The Battlefield Cyl Fog was verified (in a manner we will recall shortly). In Theorem 1.21, we present fixed-parameter asymptotics for ABMN($\kappa,\rho$) elements throughout the region $(\kappa,\rho)\in(0,1]^{2}$ in which such elements describe Nash equilibria in $\mathrm{TLP}(\kappa,\rho)$ according to Theorem 1.8. This result permits us to interrogate the validity of this premise (16) more broadly, via a two-dimensional family of models.
@@
This sequence is an important part of our apparatus for computing ABMN($\kappa,\rho$) elements. Note that $\phi_{0}=n_{0,-1}/m_{-1,0}$ equals the central ratio. The quantities $\phi_{0}$ and $\phi_{1}$ now have two meanings: as sequence elements for an ABMN($\kappa,\rho$) element, and as functions $\beta\to\phi_{i}(\kappa,\rho,\beta)$ in Definition 1.12. The coincidence is intentional, with the choice $\beta=n_{1,-1}/m_{-1,1}$ reconciling the objects, as we will see in solving the ABMN system in Section 3.
For any parameter pair $(\kappa,\rho)$ belonging to the region $W$ in (8), the orbit $\{s_{i}(x):i\in\mathbb{Z}\}$ specified in Definition 1.14 will be shown to be decreasing, for any $x\in(0,\infty)$. As we will substantiate in Section 3, $s(\phi_{i})=\phi_{i+1}$ for each $i\in\mathbb{Z}$: for any given ABMN($\kappa,\rho$) solution, $s$ acts as the unit left-shift on the just specified $\phi$-sequence. Lemma 2.7 will show that the $s$-orbit from any positive real passes exactly once through the central domain as it is next defined. As such, this lemma furnishes the existence and uniqueness claims on which the next definition depends.
@@
Here is our result offering asymptotics for each component in $(a,b,m,n)\in$ ABMN($\kappa,\rho$) in terms of distance of the index from the battlefield. The result extends the case $(\kappa,\rho)=(1,1)$ treated in [26, Theorem $2.14$]. There are three regimes in $(0,1]^{2}\setminus\{(1,1\})$: the interior, and the upper and right sides.
@@
These results suggest that the point $(1,1)$ may have singular behaviour, with the most rapid decay. This is borne out by [26, Theorem $2.14$]: $b_{-i}$ has doubly exponential leading-order decay, of the form $\exp\big\{-2\cdot 2^{i}A\}$ for some $A>0$, while $a_{-i}/b_{-i}$ also decays doubly exponentially, having the form $\exp\big\{-2^{i}A\}$ to leading order. So the premise $a_{-i}\ll b_{-i}\ll 1$ and $b_{-i}\ll a_{-i}/b_{-i}$ that we presented in (16) via the left boundary $\kappa=0^{+}$ Brownian Boost case is supported in all four regimes of $(\kappa,\rho)\in(0,1]^{2}$.

### v4

@@ 1.4. Brownian Boost and the high-noise limit
We now return to Brownian Boost. Our analysis of the game operates by regularizing it via a fine-mesh high-noise scaling of the Trail of Lost Pennies. In the first of three subsections, we advocate this discretization as a natural means of rigorous analysis of Brownian Boost. In the second, we develop the analytic framework for $\mathrm{BB}(\rho)$ equilibria and their properties that was glimpsed in Section 1.1.2. In the third, we further the presentation of our principal conclusions about $\mathrm{BB}(\rho)$ by stating Theorem 1.18, which describes equilibrium stakes and gameplay in the high-noise regime via this analytic framework.
@@ 1.4.1. The scaled high-noise trail game as a regularized Brownian Boost
As such, $\mathrm{BB}_{\kappa}(\rho)$ may be more closely compared to a scaled version $\mathrm{ScTLP}(\kappa,\rho)$ of the Trail of Lost Pennies. The scaled game operates by the rules of $\mathrm{TLP}(\kappa,\rho)$ with two changes: it is played on $\kappa\,\mathbb{Z}$, not $\mathbb{Z}$; and the running costs (1.2) that enter the net receipt formulas (1.3) now include $\kappa^{2}$-prefactors, $C_{\pm}=\kappa^{2}\sum_{k=0}^{\infty}S_{\pm}(k)$.
@@
It is natural to pose the problem of determining the ‘domain of attraction’ of discretized approximant games for $\mathrm{BB}(\rho)$. Adapting the methods of [12, 14], [19] addresses this type of question for stochastic differential games of zero-sum; while the strongly non-anticipatory framework in [8] is adapted to the non-zero-sum case. Implementing a framework such as [8]’s rigorously for $\mathrm{BB}(\rho)$ would require careful handling of the infinite horizon, non-zero-sum payoffs, and the absence of discounting in Brownian Boost. Instead we choose the concrete discretization $\mathrm{TLP}(\kappa,\rho)$ in the limit of low $\kappa$ (or high flip noise) as the rigorous point of contact with $\mathrm{BB}(\rho)$.
Now we present the ODE system, including its solutions and some important properties, that governs our characterization of $\mathrm{BB}(\rho)$ equilibria.
Here is the ODE pair that will be shown to govern time-invariant Nash equilibria in $\mathrm{BB}(\rho)$.
Definition 1.13.
$2\rho f^{1+\rho}g^{\rho}$ $=$ $\,\,\,\,\,\,\big(f^{2\rho}-g^{2\rho}\big)f\,+\,\tfrac{1}{2}f^{\prime}\big(f^{\rho}+g^{\rho}\big)^{2}\,$, (1.13)
A pair of functions $f,g:\mathbb{R}\to(0,\infty)$ is called default if $f(0)=1$ and $g(0)>0$.
In our game-theoretic interpretation of the pair $(f,g)$, the domain variable is a location on the real-line gameboard, and $f$ and g-g are spatial derivatives of expected net receipt at equilibrium for Maxine and Mina, viewed as functions of initial counter location. The minus sign attached to $g$ renders this function positive, reflecting that Mina plays to the left and improves her position by doing so.
The ODE pair arises as the coupled system of Hamilton–Jacobi–Bellman [HJB] equations associated to the non-zero-sum game $\mathrm{BB}(\rho)$. In Chapter 5, we will explain this connection with a simple but non-rigorous argument. Using Markovian balance equations and stability under momentary perturbation of stake by a given player, the argument yields necessary conditions for a stake-profile pair $(a,b):\mathbb{R}\to[0,\infty)^{2}$ to constitute a Nash equilibrium. Associated to $(a,b)$ are value functions $m,n:\mathbb{R}\to[0,\infty)$, representing the players’ expected net receipts as a function of the initial counter location. Under a differentiability hypothesis, these satisfy $m^{\prime}>0$ and $n^{\prime}<0$, and the resulting HJB conditions identify $(f,g)=(m^{\prime},-n^{\prime})$ as solutions of (1.13). (In the theory of stochastic differential games, formal derivations of HJB equations would suppose sufficient differentiability; but, in contrast to $\mathrm{BB}(\rho)$, value functions often do not enjoy that regularity, and are instead exhibited rigorously as viscosity solutions [10]: see [19] and [4] respectively for zero- and non-zero-sum treatments.)
In the preceding section, we identified explicitly the time-invariant Nash equilibria of the discrete Trail of Lost Pennies games. In the next subsection, we will state how, when $\rho\in(0,1]$, in the high-noise fine-mesh limit these discrete equilibria converge to Brownian Boost equilibria, and that the limiting equilibrium objects satisfy the ODE pair in Definition 1.13. Thus, for $\rho\in(0,1]$, the ODE system provides the analytic framework in which time-invariant Nash equilibria in $\mathrm{BB}(\rho)$ are classified.
For now, the prospect of such a characterization may provoke the question, how to solve the above pair of equations? We record the answer next, noting that currency revaluation permits us to consider only default solutions. Our analytic deductions hold whenever $\rho\in(0,\infty)$, even if the game-theoretic meaning of the $\mathrm{BB}(\rho)$ ODE pair is unsettled for $\rho>1$.
Definition 1.14.
For $\rho,\xi\in(0,\infty)$, let $S_{\rho}(\xi,\bullet):\mathbb{R}\to(0,\infty)$ denote the unique solution to the differential equation
$\frac{{\rm d}}{{\rm d}u}S_{\rho}(\xi,u)\,=\,-\frac{8\rho\,S_{\rho}(\xi,u)^{1+\rho}}{\big(1+S_{\rho}(\xi,u)^{\rho}\big)^{2}}\,,\qquad S_{\rho}(\xi,0)=\xi\,.$
Associate to this solution the pair of functions $f_{\rho}(\xi,\bullet),g_{\rho}(\xi,\bullet):\mathbb{R}\to(0,\infty)$ by means of
$f_{\rho}(\xi,u)\,=\,\exp\!\left\{2\int_{0}^{u}\Bigg(1-\frac{2}{\big(1+S_{\rho}(\xi,r)^{\rho}\big)^{2}}\,\Big(1+(1-\rho)S_{\rho}(\xi,r)^{\rho}\Big)\Bigg)\,{\rm d}r\right\}$
and
$g_{\rho}(\xi,u)\,=\,\xi\cdot\exp\!\left\{-2\int_{0}^{u}\Bigg(1-\frac{2}{\big(1+S_{\rho}(\xi,r)^{\rho}\big)^{2}}\,\Big((1-\rho)S_{\rho}(\xi,r)^{\rho}+S_{\rho}(\xi,r)^{2\rho}\Big)\Bigg)\,{\rm d}r\right\}$
for $u\in\mathbb{R}$. When $u<0$, the integrals are specified in the usual way: $\int_{a}^{b}h=-\int_{b}^{a}h$ for $a>b$.
The parameter $\xi\in(0,\infty)$ acts as an index for the space of flows $S_{\rho}(\xi,\bullet):\mathbb{R}\to(0,\infty)$. We reserve the symbol $\xi$ for this usage, calling it the flow index. This index parameterises default solutions of the $\mathrm{BB}(\rho)$ ODE pair as follows.
Theorem 1.15.
Let $\rho\in(0,\infty)$. The space of default solutions to the system (1.13) is equal to
$\Big\{\big(f_{\rho}(\xi,\bullet),g_{\rho}(\xi,\bullet)\big):\mathbb{R}\to(0,\infty)^{2}\Big\}\,$,
where the index runs over $\xi\in(0,\infty)$. For each $\xi$, we have $g_{\rho}(\xi,\bullet)=f_{\rho}(\xi,\bullet)S_{\rho}(\xi,\bullet)$.
In the discrete context, we interpreted the central ratio $n_{0,-1}/m_{-1,0}$ as a measure of local relative incentive for a turn at which the counter is at the origin. Recall that this parameter acts as the domain variable $x$ that parameterises the explicit forms of ABMN($\kappa,\rho$) elements recorded in Subsection 1.3.3. A counterpart role for Brownian Boost is played by the flow index $\xi=S_{\rho}(\xi,0)$, since this quantity is equal to the ratio $g_{\rho}(\xi,0)/f_{\rho}(\xi,0)$ of the spatial derivative (in absolute value) of expected net receipts evaluated at 0.
Given a solution $(f,g)$ of the ODE pair, the associated time-invariant stake profile $(a,b)$ is recovered explicitly.
Definition 1.16.
For $\xi\in(0,\infty)$, define functions $a_{\rho}(\xi,\bullet),b_{\rho}(\xi,\bullet):\mathbb{R}\to(0,\infty)$ by
$a_{\rho}(\xi,u)=2\rho\,\frac{f_{\rho}(\xi,u)^{1+\rho}\,g_{\rho}(\xi,u)^{\rho}}{\big(f_{\rho}(\xi,u)^{\rho}+g_{\rho}(\xi,u)^{\rho}\big)^{2}},\qquad b_{\rho}(\xi,u)=2\rho\,\frac{f_{\rho}(\xi,u)^{\rho}\,g_{\rho}(\xi,u)^{1+\rho}}{\big(f_{\rho}(\xi,u)^{\rho}+g_{\rho}(\xi,u)^{\rho}\big)^{2}}.$
The default solutions of Definition 1.13 are normalized by a derivative condition imposed at the origin, in correspondence with the scaling symmetry given by currency revaluation. For any solution arising from Definition 1.14, however, the associated global quantities
$m_{\rho}(\xi,\infty):=\int_{-\infty}^{\infty}f_{\rho}(\xi,r)\,{\rm d}r\,,\qquad n_{\rho}(\xi,-\infty):=\int_{-\infty}^{\infty}g_{\rho}(\xi,r)\,{\rm d}r$ (1.14)
are finite and positive. Interpreted as integrals of spatial derivatives of expected net receipts, these quantities represent the players’ terminal rewards in the event of victory. As such, these quantities are attributes of $\mathrm{BB}(\rho)$ game data, and, given our convention for specifying the continuum game, it is natural to use a normalization, counterpart to the discrete case in Definition 1.12, for which Maxine’s winning receipt equals one.
Definition 1.17.
A solution of the $\mathrm{BB}(\rho)$-ODE pair $(f,g):\mathbb{R}\to(0,\infty)$ is called standard if $\int_{\mathbb{R}}f(r)\,{\rm d}r=1$.
Up to trivial symmetries, the family
$\big\{(f_{\rho}(\xi,\bullet),g_{\rho}(\xi,\bullet)):\xi\in(0,\infty)\big\}$
exhausts all solutions of the $\rho$-Brownian Boost ODE pair. The switch from default to standard normalization may be effected by dilating the displayed pair by a suitable positive $\xi$-dependent factor.
Our analysis of the $\mathrm{BB}(\rho)$-ODE system will show that in fact the integral expressions specifying $m_{\rho}(\xi,\infty)$ and $n_{\rho}(\xi,-\infty)$ in (1.14) coincide for any given $\xi\in(0,\infty)$. This equality is maintained by dilation and thus holds for any normalization of solutions. Consequently, $\int_{\mathbb{R}}g(r)\,{\rm d}r=1$ for any standard solution.
@@ 1.4.3. Scaled gameplay in the high-noise trail game
Here we substantiate that for $\rho\in(0,1]$ the time-invariant equilibria of $\mathrm{BB}(\rho)$ are given by the prescription in Definitions 1.14 and 1.16, with a result showing that this description captures (in the limit of low $\kappa$) all time-invariant stake-profiles and gameplay in the scaled trail game $\mathrm{ScTLP}(\kappa,\rho)$.
@@
is the unique element of ABMN($\kappa,\rho$) with $\phi_{0}=x$ and $m_{-1,0}=\kappa$ (as well as $m_{-\infty}=n_{\infty}=0$). In view of Theorems 1.3 and  1.11, time-invariant Nash equilibria in $\mathrm{TLP}(\kappa,\rho)$ are characterized up to the trivial symmetries by these solutions, and we use them to express our result Theorem 1.18.
@@
    $\kappa^{-2}a^{\rm def}_{\lfloor\kappa^{-1}u\rfloor}(\kappa,\rho,\xi)$ $\,=\,a_{\rho}(\xi,u)\;\big(1+\ O(\kappa)\big)\,$, (1.15)
    $\kappa^{-2}b^{\rm def}_{\lfloor\kappa^{-1}u\rfloor}(\kappa,\rho,\xi)$ $\,=\,b_{\rho}(\xi,u)\;\big(1+\ O(\kappa)\big)\,$, (1.16)
    with the implicit constant in the $O$-terms being uniform in $(\rho,\xi,u)\in(0,1]\times K\times K$ for compact $K\subset(0,\infty)$.
@@
    $0,\infty)\to\mathbb{R}:t\to\kappa\,X_{\kappa,\rho}(\kappa^{-1}z,\kappa^{-2}t)\,$,
    whose domain of definition is enlarged from $\kappa^{2}\mathbb{N}$ to $0,\infty)$ by interpolation.
    Equip the space $\mathcal{C}$ of continuous functions $f:[0,\infty)\to\mathbb{R}$ with the topology of uniform convergence on compact intervals. As $\kappa\searrow 0$, this process converges weakly in $\mathcal{C}$ to the unique solution $Z:[0,\infty)\to\mathbb{R}$ of the stochastic differential equation
    ${\rm d}Z_{t}=R_{\rho}(Z_{t})\,{\rm d}t+{\rm d}W_{t},$
    with $Z_{0}=z$, where $W_{t}$ is standard Brownian motion. The function $R_{\rho}:\mathbb{R}\to(-1,1)$ is given by $R_{\rho}(u)=\frac{1-S_{\rho}(1,u)^{\rho}}{1+S_{\rho}(1,u)^{\rho}}$ or equivalently $\frac{1-S_{1}(1,\rho^{2}u)}{1+S_{1}(1,\rho^{2}u)}$. It has asymptotics
@@
Remarks: (1). In introducing the theorem, we referred to its description of $\mathrm{BB}(\rho)$ TINE as being up to trivial symmetries. To be precise, our convention for terminal receipts in $\mathrm{BB}(\rho)$ is $\lambda\cdot{\bf 1}_{E_{-}}$ for Mina and ${\bf 1}_{E_{+}}$ for Maxine. Since $m_{\rho}(\xi,\infty)$ and $n_{\rho}(\xi,-\infty)$ in (1.14) are equal for any given $\xi\in(0,\infty)$, we have $\lambda=1$ for all $\mathrm{BB}(\rho)$ TINE. The correct normalization of such TINE is the standard one, with the $\xi$-indexed TINE taking the form $\mathbb{R}\to(0,\infty)^{2}:u\to Q(\xi)^{-1}\big(a_{\rho}(\xi,u),b_{\rho}(\xi,u)\big)$, with normalization factor $Q(\xi)$ equal to the common value of $m_{\rho}(\xi,\infty)$ and $n_{\rho}(\xi,-\infty)$.
(2). The function $J(u)$ in the earlier signpost (1.1) equals $S_{1}(1,\rho^{2}u)$. The form for $J^{\prime}$ recorded there is given by taking $\rho=1$ in Definition 1.14 with a linear change of variable: see Lemma 5.1.
@@ 1.5. Robustness of inferences: the discouragement effect and asymmetric decay
1.5.1. Brownian Boost equilibrium behaviour
We present properties of the Brownian Boost equilibrium objects from Section 1.4.2 pertinent for comparing to dynamic contest theory.
Default solutions of the $\mathrm{BB}(\rho)$ ODE pair are indexed by the flow index $\xi\in(0,\infty)$, and take the form $\big(f_{\rho}(\xi,\bullet),g_{\rho}(\xi,\bullet)\big):\mathbb{R}\to(0,\infty)^{2}$ as described in Theorem 1.15. In the signposts Section 1.1.2, we indicated that the space of time-invariant Nash equilibria is invariant under spatial shifts, and is naturally indexed by a real battlefield value $v$. This translation invariance acts by shifting the spatial argument of the equilibrium profiles. The time-invariant Nash equilibria associated to default solutions are given by the stake-profile pairs $\big(a_{\rho}(\xi,\bullet),b_{\rho}(\xi,\bullet)\big)$, as specified in Definition 1.16. The flow-index and battlefield-value descriptions are reconciled in Section 5.2 where a one-to-one correspondence $(0,\infty)\longleftrightarrow\mathbb{R}$ between the two indices is presented.
When $\xi=1$, the associated battlefield value $v$ is the origin. It is enough to study this choice. The next result does so, capturing the rapid decay and the asymmetry that are characteristic of all solutions’ behaviour. Before stating it, we note that the equality $m_{\rho}(1,\infty)=n_{\rho}(1,-\infty)$ noted at the end of Section 1.4.2 has a direct game-theoretic interpretation: since these quantities represent terminal rewards, time-invariant equilibria exist in Brownian Boost only when player incentives are exactly balanced, that is, when $\lambda=1$ in the Brownian Boost setup from Section 1.1.
Proposition 1.19.
For $\rho\in(0,\infty)$, the functions $f_{\rho}(1,u)$, $g_{\rho}(1,u)$, $a_{\rho}(1,u)$ and $b_{\rho}(1,u)$ satisfy, as $u\to\infty$,
$f_{\rho}(1,u)=u^{\zeta_{f}}e^{-2u}\,\Theta(1),\qquad g_{\rho}(1,u)=u^{\zeta_{g}}e^{-2u}\,\Theta(1),$
and
$a_{\rho}(1,u)=u^{\zeta_{a}}e^{-2u}\,\Theta(1),\qquad b_{\rho}(1,u)=u^{\zeta_{b}}e^{-2u}\,\Theta(1),$
where
$\zeta_{f}=\frac{1+\rho}{2\rho^{2}},\qquad\zeta_{g}=\frac{1-\rho}{2\rho^{2}},\qquad\zeta_{a}=\frac{1+\rho}{2\rho^{2}}-1,\qquad\zeta_{b}=\frac{1-\rho}{2\rho^{2}}-1.$
As $u\to$-$\infty$, the corresponding asymptotics are
$f_{\rho}(1,u)=|u|^{\zeta_{g}}e^{-2|u|}\,\Theta(1),\qquad g_{\rho}(1,u)=|u|^{\zeta_{f}}e^{-2|u|}\,\Theta(1),$
and
$a_{\rho}(1,u)=|u|^{\zeta_{b}}e^{-2|u|}\,\Theta(1),\qquad b_{\rho}(1,u)=|u|^{\zeta_{a}}e^{-2|u|}\,\Theta(1).$
Figure $1.$ The curves $a=a_{1}(1,\bullet)$ and $f=f_{1}(1,\bullet)$, as specified by Definitions 1.14 and 1.16; $b$ and $g$ are given by reflecting in the vertical axis. Maxine’s stake profile $a$ takes maximum value $0.57$ at $u=0.25$ to two decimal places.
When $\xi=1$, players stake at unit order in a bounded neighbourhood of the origin. (As Figure 1.1 depicts in the case $\rho=1$, a player spends most as she begins to lead.) All four functions, including the stake profiles $a_{\rho}(1,\cdot)$ and $b_{\rho}(1,\cdot)$, decay rapidly, as $e^{-2|u|}$, at high distances from the origin. There is a more modest but clear asymmetry in the rate of decay, manifest in the values of the power-law exponents. When $u\gg 0$, the presumptively leading player, Maxine, is staking at normalized rate $u^{\zeta_{a}}$, above the analogous level of $u^{\zeta_{b}}$ for Mina; and this circumstance is swapped in the opposite regime $u\ll 0$. As $\zeta_{a}>\zeta_{b}$, so $\zeta_{f}>\zeta_{g}$: $1\gg m^{\prime}\gg-n^{\prime}>0$ when $u\gg 0$; integrating on $u,\infty)$, the leading player’s shortfall in expectation relative to her winning terminal receipt is seen to exceed the opponent’s excess over her losing terminal receipt. This imbalance reflects the effort of expenditure that the leading player must exert—small in an absolute sense, but large relative to the opponent’s—in order to convert a likely victory.
@@ 1.5.2. Fixed-parameter ABMN($\kappa,\rho$) asymptotics, and asymmetric decay
These themes are apparent in $\mathrm{BB}(\rho)$ from Proposition 1.19, wherein the flow index choice $\xi=1$ corresponds to the battlefield value $v$ being the origin. Stake-profiles $\Theta(1)$ in a compact neighbourhood containing 0, while satisfying
$a_{-u}\ll b_{-u}\ll a_{-u}/b_{-u}\ll 1\,\,\,\,\textrm{for}\,\,\,\,u\gg 0:$ (1.17)
in negative territory, where Mina leads, stakes have fallen exponentially, the more so for Maxine, though the decay in stake ratio has a more modest polynomial rate. ‘Battlefield Cyl Fog’ (cut your losses, foot on gas) is a mnemonic for the premise (1.17), the phrases descriptive of the trailing and leading player’s respective approach far from the battlefield.
In [24], TLP(1,1)(1,1) was studied: the Trail of Lost Pennies without flip moves whose turn outcomes are decided by the simple $\frac{a}{a+b}$ ‘lottery’ rule. The Battlefield Cyl Fog was verified (in a manner we will recall shortly). In Theorem 1.22, we present fixed-parameter asymptotics for ABMN($\kappa,\rho$) elements throughout the region $(\kappa,\rho)\in(0,1]^{2}$ in which such elements describe Nash equilibria in $\mathrm{TLP}(\kappa,\rho)$ according to Theorem 1.3. This result permits us to interrogate the validity of this premise (1.17) more broadly, via a two-dimensional family of models.
@@
This sequence is an important part of our apparatus for computing ABMN($\kappa,\rho$) elements. Note that $\phi_{0}=n_{0,-1}/m_{-1,0}$ equals the central ratio. Being a measure of single-turn relative incentive for play at the origin, it is a discrete counterpart to the flow index $\xi$ in Brownian Boost.
Note that the quantities $\phi_{0}$ and $\phi_{1}$ now have two meanings: as sequence elements for an ABMN($\kappa,\rho$) element, and as functions $\beta\to\phi_{i}(\kappa,\rho,\beta)$ in Definition 1.7. The coincidence is intentional, with the choice $\beta=n_{1,-1}/m_{-1,1}$ reconciling the objects, as we will see in solving the ABMN system in Chapter 3. The meaning of $\phi_{0}$ and $\phi_{1}$ will, we hope, be clear from context: specified by the formulas when $\beta\in(0,\infty)$ is a free parameter, and as sequence elements when an ABMN($\kappa,\rho$) element is given.
For any parameter pair $(\kappa,\rho)$ belonging to the region $W$ in (1.7), the orbit $\{s_{i}(x):i\in\mathbb{Z}\}$ specified in Definition 1.9 will be shown to be decreasing, for any $x\in(0,\infty)$. As we will substantiate in Chapter 3, $s(\phi_{i})=\phi_{i+1}$ for each $i\in\mathbb{Z}$: for any given ABMN($\kappa,\rho$) solution, $s$ acts as the unit left-shift on the just specified $\phi$-sequence. Lemma 2.7 will show that the $s$-orbit from any positive real passes exactly once through the central domain as it is next defined. As such, this lemma furnishes the existence and uniqueness claims on which the next definition depends.
@@
Here is our result offering asymptotics for each component in $(a,b,m,n)\in$ ABMN($\kappa,\rho$) in terms of distance of the index from the battlefield. The result extends the case $(\kappa,\rho)=(1,1)$ treated in [24, Theorem $2.14$]. There are three regimes in $(0,1]^{2}\setminus\{(1,1\})$: the interior, and the upper and right sides.
@@
These results suggest that the point $(1,1)$ may have singular behaviour, with the most rapid decay. This is borne out by [24, Theorem $2.14$]: $b_{-i}$ has doubly exponential leading-order decay, of the form $\exp\big\{-2\cdot 2^{i}A\}$ for some $A>0$, while $a_{-i}/b_{-i}$ also decays doubly exponentially, having the form $\exp\big\{-2^{i}A\}$ to leading order. So the premise $a_{-i}\ll b_{-i}\ll 1$ and $b_{-i}\ll a_{-i}/b_{-i}$ that we presented in (1.17) via the left boundary $\kappa=0^{+}$ Brownian Boost case is supported in all four regimes of $(\kappa,\rho)\in(0,1]^{2}$.